I am trying to test an equation for symmetry. The equation is x^2 - y^3 + 4.
I am starting by trying to find the intercepts. I think I found the intercepts for x - (2,0), (-2,0). I am having trouble finding the intercepts for y.
I substitute 0 for X and get 0 + y^3 + 4. I convert that to y^3 = -4. I am not sure where to go from there. Can I get a cube root of four? Solving for y, would my answer be the cube root of four i? Does that mean there is no y intercept?
I've confused myself... Thanks for any help.

Answer by stanbon(75887)

I am trying to test an equation for symmetry. The equation is x^2 - y^3 = 4.
------
y-axis symmetry if you substitute -x for x
and it would give you the same equation.
Example: y = x^2
----------------
x-axis symmetry if you substitute -y for y
and it would give you the same equation.
Example: x = y^2
----------------------------------------
origin symmetry if you substitute -x for x
and -y for y and get the same equation.
Example: y = x^3
======================
Cheers,
Stan H.
======================